OpenStudy (anonymous):

Graph the equations using the x and y intercepts. (x/4)-(y/5)=1 I am stuck on this and I cannot figure out how to get this into the right equation/form to be able to graph it.

OpenStudy (callisto):

You may start with finding the x and y intercepts first. To find the x-intercept, you simply put y=0 into the given equation and solve x. The x value you find is the x-intercept. Similarly, you can find the y-intercept by putting x=0 into the given equation and solve y. The y value you find is the y-intercept. After that, you can draw a coordinate plane. Mark the two intercepts you've found and join the two points with a line. That would be the graph for the equation. You are also reminded to draw the equation next to the line you've drawn.

OpenStudy (anonymous):

so I would do this: (x/4)-(0/5)=1 when I subtract the two and find a common denominator I get (5x/20) this is where I am confused with the fractions I do not know what to do next. If it were not a fraction to begin with I would be ok.

OpenStudy (callisto):

Alright. We'll go through it together. We're looking for the x-intercept, so we put y=0 into the give equation. (x/4)-(0/5)=1. Now, what is 0/5?

OpenStudy (anonymous):

It is 0. I have an ordered pair of (8,5) so far.

OpenStudy (callisto):

It is 0, yes. How did you get the ordered pair (8,5)?

OpenStudy (anonymous):

if x=8, and y=5 (8/4)-(5/5)=1 2 - 1 =1 also if you do it the long way and find common denominator it will still work out the same. Also (0,5) works. I believe this to be a straight line on the graph

OpenStudy (anonymous):

Never mind the second part about (0,5) and the straight line, not sure what I was thinking there,

OpenStudy (anonymous):

Here's one (4,0)

OpenStudy (callisto):

(0,5) works because it's a point on the line. If you substitute two numbers (x,y) into the equation ad you've got both sides equal, then it's a point on the line.

OpenStudy (callisto):

(4,0) is right. How did you get that?

OpenStudy (anonymous):

I basically used the same concept and I just started plugging in numbers for x and y that would fit into the equation

OpenStudy (anonymous):

But, if (4,0) is right (0,5) cannot be

OpenStudy (callisto):

Why can't (0,5) be right?

OpenStudy (anonymous):

My first point is (8,5) if I also have a point (4,0) then I cannot have one that is (0,5), however (0,-5) fits into the equation and on the line

OpenStudy (callisto):

Oh! I read something wrong. Let's see... For the point (0,5) Left side = 0/4 - (5/5) = -1 Right side = 1 \(\ne\) left side So, (0,5) is not a point on the line. For the point (0,-5) Left side = 0/4 - (-5)/5 = 1 Right side = 1 = left side. So, (0,-5) is a point on the line.

OpenStudy (anonymous):

Awesome I got it!!! Thank you for working with me on this!!!

OpenStudy (callisto):

You're welcome. But do you know what the x-intercept and y-intercept are?

OpenStudy (anonymous):

(4,-6)

OpenStudy (anonymous):

Sorry it took me so long, I am at work and I had to do a safety check.

OpenStudy (callisto):

No.... 

OpenStudy (anonymous):

Yes that is the line I got also

OpenStudy (callisto):